### v1

# GeoBEV: Learning Geometric BEV Representation for Multi-view 3D Object Detection

Abstract
Bird’s-Eye-View (BEV) representation has emerged as a mainstream paradigm for multi-view 3D object detection, demonstrating impressive perceptual capabilities. However, existing methods overlook the geometric quality of BEV representation, leaving it in a low-resolution state and failing to restore the authentic geometric information of the scene. In this paper, we identify the reasons why previous approaches are constrained by low BEV representation resolution and propose Radial-Cartesian BEV Sampling (RC-Sampling), enabling efficient generation of high-resolution dense BEV representations without the need for complex operators. Additionally, we design a novel In-Box Label to substitute the traditional depth label generated from the LiDAR points. This label reflects the actual geometric structure of objects rather than just their surfaces, injecting real-world geometric information into the BEV representation. Furthermore, in conjunction with the In-Box Label, a Centroid-Aware Inner Loss (CAI Loss) is developed to capture the fine-grained inner geometric structure of objects. Finally, we integrate the aforementioned modules into a novel multi-view 3D object detection framework, dubbed GeoBEV. Extensive experiments on the nuScenes dataset exhibit that GeoBEV achieves state-of-the-art performance, highlighting its effectiveness.
1 Introduction
Multi-view 3D object detection stands as a prominent perception paradigm for cost-effective autonomous driving. Presently, many camera-only detectors (Huang et al. 2021; Huang and Huang 2022a; Li et al. 2023c, 2022, b; Yang et al. 2023) transform image features into Bird’s-Eye-View (BEV) space and directly perform detection on the BEV features, demonstrating competitive performance. This illustrates the substantial advantages of BEV representation in preserving comprehensive scene information, making it more adept for vision-centric autonomous driving perception than isolated image features in perspective space (Park et al. 2021; Wang et al. 2021, 2022b, 2022a).
Figure 1: Comparison between BEV representations. BEVDepth (Li et al. 2023c) is chosen as the baseline. Larger BEV size, RC-Sampling and In-Box Label are gradually applied on the baseline. The boxes represent the ground truth of the scene and brightness in BEV representation reveals the norm of the features. The background is filtered out in (b)-(e) to show the difference in the foreground.
As the cornerstone of BEV-based approaches, the BEV representation embodies both contextual semantic information and depth geometric information. The former is derived from transformed image features, while the latter originates from the correlation between image features and BEV features, calculated using camera parameters. Both types of information are indispensable for precise 3D object detection. However, the geometric quality of BEV representation has never received sufficient attention, and the limitations of low-resolution representations always arise: (1) For LSS-based methods that transform pseudo-points into BEV representation (Xie et al. 2022; Philion and Fidler 2020; Huang et al. 2021; Huang and Huang 2022a), the density imbalance of pseudo-points results in many positions within the BEV representation having vacant features, as depicted in the peripheral areas of Fig. 1(a). Furthermore, the sparsity will further increase along with the resolution of the BEV representation, as shown in Fig. 1(c). (2) For Transformer-based methods that employ cross-attention to retrieve image features (Li et al. 2022; Yang et al. 2023; Jiang et al. 2023; Li et al. 2023a), the elevated resolution of the BEV representation leads to a rapid escalation in computational costs. The lack of explicit depth distribution also limits their ability to restore accurate geometric information. Overall, regardless of the type of method, their respective shortcomings restrict the generation of high-resolution BEV representations, thus compromising the ability to accurately depict objects.
To solve the drawbacks of the existing feature transformation mechanisms, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to efficiently generate dense BEV representation with high resolution. Initially, Radial BEV features are obtained through high-dimensional matrix multiplication between the transposed image features and depth scores. Subsequently, bilinear sampling is employed to retrieve the corresponding Radial BEV features for populating the BEV features in Cartesian coordinates. The sampling procedure mitigates the sparsity of the BEV representation and incurs minimal computational cost even with increased resolution. As shown in Fig. 1(d), applying RC-Sampling effectively avoids the vacant features that appear in Fig. 1(c), enhancing the geometric quality of BEV representations. RC-Sampling outperforms state-of-the-art feature transformation approaches on both precision and efficiency without using complex operators.
Enhancing the accuracy of geometric information is equally crucial for BEV-based approaches. Some methods attempt to supervise predicted depth scores by utilizing the depth values of LiDAR points as depth labels (Reading et al. 2021; Li et al. 2023c, b; Wang et al. 2022c; Zhang et al. 2023a). However, the LiDAR labels only record the depth of object surfaces near the camera, failing to represent the actual geometric structure of objects in the real-world space. We propose In-Box Label that can significantly enhance the geometric quality of BEV representations without adopting extra modules. We first check whether the generated pseudo-points are within the GT boxes to obtain binary labels. These labels, called vanilla In-Box Label, can effectively incentivize the network to assign high depth scores to where the objects are actually located. Nonetheless, they may lead to feature confusion caused by object occlusion or background pseudo-points being wrongly boxed. We ameliorate those issues to enhance its accuracy in reflecting the geometric structure of the scene. In conjunction with the utilization of In-Box Labels, Centroid-Aware Inner Loss is also proposed to capture the fine-grained inner geometric structure of objects. After applying the In-Box Label, the authentic geometric structures of objects are clearly presented as shown in Fig. 1(e), and more precise detection is facilitated.
We integrate the aforementioned modules into a novel multi-view 3D object detector, dubbed GeoBEV, and carry out extensive experiments on the nuScenes dataset. The major contributions of this paper can be summarized as:
  • We propose Radial-Cartesian BEV Sampling to conveniently acquire Cartesian BEV features by bilinearly sampling Radial BEV features, which enables the efficient generation of high-resolution dense BEV representations, facilitating the recovery of fine-grained geometric details within the scene.
  • We design a novel In-Box Label to supervise the predicted depth score, which better reflects the actual geometric structure of the object than the LiDAR label, injecting authentic geometric information into the BEV representation.
  • Extensive experiments are conducted on the nuScenes Dataset, and GeoBEV reaches newly state-of-the-art results for multi-view 3D object detection, highlighting its effectiveness.
2 Related Work
2.1 Depth Prediction Based BEV Representation
Due to the inherent limitations of the image modality in providing the indispensable explicit depth information required for 3D object detection, predicting the depth distribution of image elements becomes a natural choice. Early methods like OFT (Roddick, Kendall, and Cipolla 2018) assume that the depth distribution of image elements is uniform and all voxels along the ray starting from the camera share the same features. Lately, LSS (Philion and Fidler 2020) enables networks to adaptively predict depth distributions and weight image features to generate pseudo-points at corresponding depth values, which are then accumulated into BEV representations. BEVDet (Philion and Fidler 2020) employs LSS to construct the detection framework and proposes a data augmentation strategy on BEV features. BEVDet4D (Huang and Huang 2022a) integrates the BEV features from past frames to assist in predicting the velocity of the objects in the current frame.
In order to obtain more accurate depth information, CaDDN (Reading et al. 2021) projects the LiDAR points onto the image to provide supervision on the predicted depth distribution. BEVDepth (Li et al. 2023c) also considers the influence of the camera’s internal and external parameters and further optimizes the depth distribution after the supervised prediction. BEVStereo (Li et al. 2023b) utilizes the structure of multi-view stereo to obtain more reliable depth distributions and performs some optimizations to minimize memory usage. TiG-BEV (Huang et al. 2022) sets several key points to learn the local depth structure of the scene better. SA-BEV (Zhang et al. 2023a) segments the images to get the foreground-only BEV features and improves depth distribution via multi-task learning. BEV-IO (Zhang et al. 2023b) adopts instance occupancy prediction modules as a complement to depth prediction. FB-BEV (Li et al. 2023d) combines forward and backward projection, compensating for the deficiencies in both existing methods. However, these attempts to optimize depth prediction fail to avoid geometrical fallacies. In this paper, we propose In-Box Labels that contain comprehensive geometric information, significantly enhancing the geometric perception of BEV representation.
Figure 2: Overall architecture of GeoBEV. The multi-view images are processed to provide the image features and depth scores. The depth scores are supervised by the In-Box Label that restores authentic geometric structures of objects through the Centroid-Aware Inner Loss. Radial-Cartesian BEV Sampling then conveniently generates dense BEV representation with high resolution.
2.2 Transformer Based BEV Representation
With the help of the attention mechanism, Transformer-based detectors can adaptively retrieve and combine image features to obtain dense BEV representation. BEVFormer (Li et al. 2022) uses deformable attention to find the corresponding image features, and additionally fuses the BEV representation from previous frames. BEVFormerV2 (Yang et al. 2023) introduces a detection head in perspective view to make the image features that will be retrieved more suitable for 3D detection. PolarFormer (Jiang et al. 2023) argues that polar coordinates are more competent for ego car perception and designs a polar detection head to deal with the irregular polar grids. DFA3D (Li et al. 2023a) utilizes the explicit depth distribution when doing the cross attention from image to BEV and simplifies the 3D Transformer into the 2D Transformer equivalently.
Several Transformer-based detectors regard the objects as queries and do not generate explicit BEV representation. DETR3D (Wang et al. 2022b) follows DETR (Carion et al. 2020; Zhu et al. 2020) series detectors and interacts object queries with multi-view image features to complete the object information for detection. PETR (Liu et al. 2022) embeds 3D position information into the image features, making the object queries easier to obtain spatial information. PETRv2 (Liu et al. 2023a) extends PETR for temporal modeling and adds map queries for other perception tasks. StreamPETR (Wang et al. 2023) propagates long-term historical information and achieves comparable accuracy to classical LiDAR based detectors. Sparse4D (Lin et al. 2022) assigns multiple 4D key points to aggregate multi-view/scale/timestamp image features. Sparse4Dv2 (Lin et al. 2023) uses the recurrent method to realize the transmission of temporal information, avoiding multiple sampling to improve efficiency. Nevertheless, the existing Transformer-based BEV representation methods involve complex computation processes and high latency, making it difficult to generalize to high-resolution BEV feature situations. In this paper, we propose Radial-Cartesian BEV Sampling to efficiently generate high-resolution dense BEV representations, facilitating the recovery of fine-grained geometric details within the scene.
3 Method
3.1 Overall Architecture
The overall architecture of our proposed GeoBEV is shown in Fig. 2. Firstly, the multi-view images are processed to provide the image features that will be transformed into BEV features and their corresponding depth scores. Then the In-Box Label is created and utilized to supervise the depth scores to restore the actual distribution of the objects in BEV space. The Centroid-Aware Inner Loss is adopted during the depth supervision to let the model learn the inner structure of the objects. Finally, Radial-Cartesian BEV Sampling is proposed to conveniently generate dense BEV representation with high resolution, which effectively saves the fine-grained geometric information of the scene.
Figure 3: The illustration of Radial-Cartesian BEV Sampling. After high-dimensional matrix multiplication between the transposed image features and depth scores, the $H$ dimension is squeezed and the radial BEV features can be directly obtained.
3.2 Radial-Cartesian BEV Sampling
Increasing the resolution of BEV features can significantly enhance its detail and make it easier for perception modules to extract fine-grained geometric information. However, the drawbacks of current methods limit the resolution of BEV representations. For LSS-based methods (Li et al. 2023c, b; Zhang et al. 2023a; Li et al. 2023d), the density imbalance of pseudo-points makes the signal strength of BEV features decays from the center to the periphery, and some places even have the vacant features. The sparsity will be further increased for BEV features in high resolution. FB-BEV (Li et al. 2023d) applies backward projection to fill these vacant features but relies on precisely predicting RoI from the sparse BEV features. Transformer-based methods(Li et al. 2022; Yang et al. 2023; Jiang et al. 2023; Qin et al. 2023) can create dense BEV features, but the computational cost increases rapidly when the resolution of BEV features increases.
Here, we propose Radial-Cartesian BEV Sampling (RC-Sampling) to conveniently generate dense BEV features with high resolution. The radial BEV features are first obtained by high-dimensional matrix multiplication between transposed image features and the predicted depth scores as shown in Fig. 3. Regardless of batch size, the image features can be denoted as $\mathbb{I}\in\mathbb{R}^{C\times H\times W}$, where $C,H,W$ represent the channel number, height and weight of the image features. The depth scores can be denoted as $\mathbb{D}\in\mathbb{R}^{D\times H\times W}$, where $D$ represents the number of discrete depth values. To get the radial BEV features, the $H$ dimension should be squeezed, so that the image features and depth scores are transposed to $\hat{\mathbb{I}}\in\mathbb{R}^{W\times C\times H}$ and $\hat{\mathbb{D}}\in\mathbb{R}^{W\times H\times D}$ respectively. The high-dimensional matrix multiplication is then applied to the last two dimensions of $\hat{\mathbb{I}}$ and $\hat{\mathbb{D}}$ and a new matrix is created by:
$\hat{\mathbb{B}}_{R}=\hat{\mathbb{I}}\hat{\mathbb{D}},$ (1)
where $\hat{\mathbb{B}}_{R}\in\mathbb{R}^{W\times C\times D}$. After transposing $\hat{\mathbb{B}}_{R}$ into $\mathbb{B}_{R}\in\mathbb{R}^{C\times D\times W}$, the radial BEV features are obtained.
The $B_{R}$ needs to be transformed into Cartesian coordinates for subsequent detection. We pre-define the coordinates of Cartesian BEV features $\mathbb{B}_{C}\in\mathbb{R}^{C\times X\times Y}$, where $X,Y$ denote the required resolution, and simply project them on the $\mathbb{B}_{R}$. The bilinear sampling is utilized to retrieve the corresponding features, which can be represented by :
$\mathbb{B}_{C}(x,y)={\rm BilinearSample}(\mathbb{B}_{R},{\rm Project}(x,y)),$ (2)
where ${\rm Project}(x,y)$ denotes the coordinates of the projected point $(x,y)$ on $\mathbb{B}_{R}$. Using bilinear sampling instead of pooling the sparse pseudo-points guarantees that each position in the BEV representation has valid information. The sampled $\mathbb{B}_{C}$ can represent the geometric information of the scenes well because the explicit depth distribution has already been incorporated into the $\mathbb{B}_{R}$. When $\mathbb{B}_{C}$ with high resolution are used for perception, the $\mathbb{I}$ and $\mathbb{D}$ with high resolution are also required to provide fine-grained geometric information of the scene. In practice, several convolutions are utilized to increase $H$ and $W$ of $\mathbb{I}$ and $\mathbb{D}$.
Compared with other feature transformation methods, RC-Sampling does not require the generation of memory-expensive 4D frustum features (Reading et al. 2021; Philion and Fidler 2020; Huang et al. 2021), the utilization of deployment-unfriendly custom operators (Li et al. 2023c; Huang and Huang 2022b; Liu et al. 2023b) or the computation-expensive cross-attention mechanism (Li et al. 2022; Yang et al. 2023; Li et al. 2023a). It is entirely the combination of the most basic official operators. RC-Sampling not only shows the advantage of generating high-resolution BEV representation, but also outperforms the state-of-the-art feature transformation methods, such as BEVPoolv2 (Huang and Huang 2022b) and DFA3D (Li et al. 2023a), on both precision and efficiency.
(a) LiDAR Label
(b) Vanilla Label
(c) Occlusion
(d) Background
(e) Behind Surface
(f) CAI Loss
Figure 4: Illustration of the associated design of In-Box Label. $H$, $W$ and $D$ represent the height, width and channel number of depth scores. The boxes are the GT boxes. The red squares and black dots denote the positive and negative classes of the In-Box Label. The triangles are the LiDAR points. The crosses in blue represent the pseudo-points that are not supervised. The deeper color in the last figure means higher loss weight.
3.3 In-Box Label
Current LSS-based detectors (Li et al. 2023c, b, d; Zhang et al. 2023a) utilize the depth of LiDAR points as the label to supervise the depth score of each pseudo-point and thus attach the geometric information of the scene to the BEV representation. However, as shown in Fig. 4(a), the LiDAR label only records the depth of the object surfaces closer to the camera, instead of the actual geometric structure of the objects. The lack of objects’ complete geometric information hinders the subsequent BEV encoder and detection head from precisely recognizing their size and orientation. Besides, the LiDAR points do not exactly match pseudo-points in space and will introduce errors that can not be eliminated. To overcome the drawbacks of the LiDAR label, we propose the In-Box Label that can be easily obtained from the 3D coordinate of pseudo-points and the GT boxes.
Denote the 3D coordinate of a pseudo-point generated from image features as $p\in\mathbb{R}^{3}$ and the space within a GT box as $B$, then the vanilla In-Box Label of $p$ can be formulated as:
$L_{inbox}=\begin{cases}1,&p\in\bigcup\limits_{i=1}^{N}B_{i}\\ 0,&p\notin\bigcup\limits_{i=1}^{N}B_{i}\end{cases}$ (3)
where $N$ is the total number of GT boxes. It means $p$ is regarded as positive if it is within any GT boxes. From Fig. 4(b), it can be found that the In-Box Label describes the actual geometric structure of objects well and can lead the network to fill the GT boxes with valid features in BEV space as shown in Fig. 1(e).
Although the vanilla In-Box Label reflects the actual geometric structure of objects well, it may cause mismatches between image features and BEV representation of objects. For instance, since Object A in Fig. 4(c) is closer to the camera than Object B, the image records the information of Object A when occlusion occurred (occlusion region is shown by red dotted box). If the blue crosses are regarded as positive, the network will be encouraged to give a high depth score there and mix the information of Object A and Object B, which is harmful for precise detection. A better way is not to supervise the pseudo-points within the occluded region and let the network learn to give a proper depth score by itself. A similar situation happens when objects have irregular shapes, as shown in Fig. 4(d). Not all pseudo-points within the box record the information object and they should also be ignored during training. We use the HTC (Chen et al. 2019) pre-trained on nuImages (Caesar et al. 2020) to provide the mask of objects and filter out the background pseudo-points within GT boxes.
As for the background regions where no GT boxes are available, the LiDAR label is still utilized to supervise the depth scores. It is to make the network learn the whole depth distribution of the scene and locate objects more precisely. Since the LiDAR label reflects the depth of the surfaces while In-Box Label records the actual spatial distribution, we modify the LiDAR label to resolve this divergence between the optimization direction of foreground and background. As shown in Fig. 4(e), we also ignore the pseudo-points behind the background surface, which are used to be regarded as negative. It lets the network adaptively predict how “thick” is the ground and the surrounding buildings and also alleviates the imbalance between the scale of positive and negative.
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
BEVDet (Huang et al. 2021) ResNet50 256$\times$704 1 0.298 0.379 0.725 0.279 0.589 0.860 0.245
BEVDet4D (Huang and Huang 2022a) ResNet50 256$\times$704 2 0.322 0.457 0.703 0.278 0.495 0.354 0.206
PETRv2 (Liu et al. 2023a) ResNet50 256$\times$704 2 0.349 0.456 0.700 0.275 0.580 0.437 0.187
BEVDepth (Li et al. 2023c) ResNet50 256$\times$704 2 0.351 0.475 0.639 0.267 0.479 0.428 0.198
BEVStereo (Li et al. 2023b) ResNet50 256$\times$704 2 0.372 0.500 0.598 0.270 0.438 0.367 0.190
STS (Wang et al. 2022c) ResNet50 256$\times$704 2 0.377 0.489 0.601 0.275 0.450 0.446 0.212
FB-BEV (Li et al. 2023d) ResNet50 256$\times$704 3 0.378 0.498 0.620 0.273 0.444 0.374 0.200
SA-BEV (Zhang et al. 2023a) ResNet50 256$\times$704 2 0.387 0.512 0.613 0.266 0.352 0.382 0.199
GeoBEV ResNet50 256$\times$704 2 0.415 0.535 0.533 0.265 0.419 0.298 0.214
DETR3D (Wang et al. 2022b) ResNet101 900$\times$1600 1 0.349 0.434 0.716 0.268 0.379 0.842 0.200
PETR (Liu et al. 2022) ResNet101 512$\times$1408 1 0.366 0.441 0.717 0.267 0.412 0.834 0.190
PETRv2 (Liu et al. 2023a) ResNet101 900$\times$ 1600 2 0.421 0.524 0.681 0.267 0.357 0.377 0.186
BEVDepth (Li et al. 2023c) ResNet101 512$\times$1408 2 0.412 0.535 0.565 0.266 0.358 0.331 0.190
CAPE (Xiong et al. 2023) ResNet101 512$\times$1408 2 0.431 0.533 - - - - -
STS (Wang et al. 2022c) ResNet101 512$\times$1408 2 0.431 0.542 0.525 0.262 0.380 0.369 0.204
SA-BEV (Zhang et al. 2023a) ResNet101 512$\times$1408 2 0.441 0.549 - - - - -
GeoBEV ResNet101 512$\times$1408 2 0.479 0.582 0.498 0.254 0.335 0.285 0.204
Table 1: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes val set. means the interval between frames may be longer than 0.5s.
3.4 Centroid-Aware Inner Loss
When replacing the LiDAR label with the In-Box Label, the former depth loss should also be changed. We propose Centroid-Aware Inner Loss (CAI Loss) to cooperate with the characteristics of the In-Box Label, which also encourages the model to learn the inner structure of the objects.
First of all, the activation function of the depth scores is changed. The former methods use softmax as the activation function to match the one-hot LiDAR label. It centralizes the depth score on one specific value by:
$\hat{d}_{i}=\frac{e^{d_{i}}}{\sum_{j=1}^{D}e^{d_{j}}},$ (4)
where $d$ is the raw depth score and $\hat{d}$ is the activated one. When the In-Box Label is utilized, the network should give all the pseudo-points within the GT boxes high depth scores. We choose sigmoid as the activation function to independently normalize the depth scores of each pseudo-point within $[0,1]$ by:
$\hat{d}_{i}=\frac{1}{1+e^{-d_{j}}}.$ (5)
The type of depth loss is also changed. When using the LiDAR label to supervise the depth scores, the discrete depth values are regarded as different classes and the depth loss is the mean cross entropy loss of every element of image features. But when using the In-Box Label, it turns into a binary classification of every pseudo-points and severe category imbalance occurs. As a result, focal loss (Lin et al. 2017) is adopted to balance the loss of the different classes to achieve better accuracy. The loss can be calculated by:
$\mathcal{L}_{focal}(p,y)=\begin{cases}-(1-\alpha)p^{\gamma}\log(1-p),&y=0\\ -\alpha(1-p)^{\gamma}\log(p),&y=1\end{cases},$ (6)
where $y$ and $p$ represent the label and the predicted score, $\alpha$ and $\gamma$ are the adjustable parameters.
Besides the spatial distribution of objects, their inner geometric information can also benefit the detection result. Inspired by Centroid-Aware Sampling (Zhang et al. 2022), we give positive pseudo-points different loss weights according to their relative position in the GT boxes. The Centroid-Aware Inner Weight is calculated by:
$W_{CAI}=\sqrt[3]{\frac{\min(f,b)}{\max(f,b)}\times\frac{\min(l,r)}{\max(l,r)}\times\frac{\min(u,d)}{\max(u,d)}},$ (7)
where $f,b,l,r,u,d$ represent the distance of a pseudo-point to the front, back, left, right, up and down surfaces of the GT box, respectively. Only the weight of positive pseudo-points needs to be calculated and the pseudo-point closer to the centroid of an object will have a higher weight as shown in Fig. 4(f). The weights are directly multiplied over the focal loss of positive pseudo-points and the CAI Loss is calculated by:
$\mathcal{L}_{CAI}(p,y)=\begin{cases}-(1-\alpha)p^{\gamma}\log(1-p),&y=0\\ -W_{CAI}\alpha(1-p)^{\gamma}\log(p),&y=1\end{cases}.$ (8)
Using the weighted loss will encourage the pseudo-points near the object centroids to have higher depth scores than the ones near the GT box surfaces. The inner geometric information is thus represented by the depth score differences.
4 Experiments
4.1 Dataset and Metrics
We evaluate our proposed method on the nuScenes (Caesar et al. 2020) dataset, a commonly used large-scale autonomous driving benchmark. It contains diverse 1000 scenarios collected from the real world, each lasting for around 20 seconds. The key samples are annotated at 2Hz and each sample is provided with the data collected from six cameras around the car, one LiDAR on the top of the car and five radars. The 1000 scenarios are split into training set (750 scenarios), validation (150 scenarios) and test set (150 scenarios). The main metric of the nuScenes dataset for 3D object detection is the nuScenes Detection Score (NDS). Except for the commonly used mean average precision (mAP), NDS is also related to five metrics that only take true positive objects into account, including mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), mean Average Attribute Error (mAAE).
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
FCOS3D (Wang et al. 2021) ResNet101 900$\times$1600 1 0.358 0.428 0.690 0.249 0.452 1.434 0.124
DETR3D (Wang et al. 2022b) VoVNet-99 900$\times$1600 1 0.412 0.479 0.641 0.255 0.394 0.845 0.133
BEVDet (Huang et al. 2021) Swin-B 900$\times$1600 1 0.424 0.488 0.524 0.242 0.373 0.950 0.148
PETR (Liu et al. 2022) VoVNet-99 900$\times$1600 1 0.441 0.504 0.593 0.249 0.383 0.808 0.132
BEVFormer (Li et al. 2022) VoVNet-99 900$\times$1600 4 0.481 0.569 0.582 0.256 0.375 0.378 0.126
BEVDet4D (Huang and Huang 2022a) Swin-B 640$\times$1600 2 0.451 0.569 0.511 0.241 0.386 0.301 0.121
PolarFormer (Jiang et al. 2023) VoVNet-99 900$\times$1600 2 0.493 0.572 0.556 0.256 0.364 0.440 0.127
PETRv2 (Liu et al. 2023a) VoVNet-99 640$\times$1600 2 0.490 0.582 0.561 0.243 0.361 0.343 0.120
BEVDepth (Li et al. 2023c) VoVNet-99 640$\times$1600 2 0.503 0.600 0.445 0.245 0.378 0.320 0.126
BEVStereo (Li et al. 2023b) VoVNet-99 640$\times$1600 2 0.525 0.610 0.431 0.246 0.358 0.357 0.138
CAPE (Xiong et al. 2023) VoVNet-99 640$\times$1600 2 0.525 0.610 0.503 0.242 0.361 0.306 0.114
FB-BEV (Li et al. 2023d) VoVNet-99 640$\times$1600 10 0.537 0.624 0.439 0.250 0.358 0.270 0.128
SA-BEV (Zhang et al. 2023a) VoVNet-99 640$\times$1600 2 0.533 0.624 0.430 0.241 0.338 0.282 0.139
GeoBEV VoVNet-99 640$\times$1600 2 0.543 0.635 0.409 0.234 0.317 0.284 0.122
Table 2: Comparison with previous state-of-the-art multi-view 3D detectors on the nuScenes test set. means the interval between frames may be longer than 0.5s.
Method Backbone Image Size Frames mAP NDS mATE mASE mAOE mAVE mAAE
BEVFormerv2 (Yang et al. 2023) ResNet50 - - 0.423 0.529 0.618 0.273 0.413 0.333 0.188
SOLOFusion (Park et al. 2022) ResNet50 256$\times$704 17 0.427 0.534 0.567 0.274 0.511 0.252 0.181
BEVPoolv2 (Huang and Huang 2022b) ResNet50 256$\times$704 9 0.406 0.526 0.572 0.275 0.463 0.275 0.188
VideoBEV (Han et al. 2023) ResNet50 256$\times$704 8 0.422 0.535 0.564 0.276 0.440 0.286 0.198
Sparse4Dv2 (Lin et al. 2023) ResNet50 256$\times$704 - 0.439 0.539 0.598 0.270 0.475 0.282 0.179
StreamPETR (Wang et al. 2023) ResNet50 256$\times$704 8 0.450 0.550 0.613 0.267 0.413 0.265 0.196
GeoBEV ResNet50 256$\times$704 8 0.479 0.575 0.496 0.261 0.438 0.236 0.216
Table 3: Comparison with the long-term temporal methods on the nuScenes val set. Benefited from the perspective-view pre-training.
4.2 Implementation Details
We adopt the BEVDepth (Li et al. 2023c) as the baseline and add our proposed In-Box Label and Radial-Cartesian BEV Sampling to build GeoBEV. When comparing with state-of-the-art methods, we follow the commonly used configurations. For the experiments on the nuScenes validation set, the ResNet50 and ResNet101 are adopted as the backbone to process the multi-view image downsampled to 256$\times$704 and 512$\times$1408, respectively. When evaluating on the nuScenes test set, the VoVNet-99 (Lee et al. 2019) pre-trained by DD3D (Park et al. 2021) is adopted as the backbone to process the images cropped to 640$\times$1600. These models are trained for 20 epochs with CBGS strategy (Zhu et al. 2019). Except for regular data augmentation in the image and BEV space, the BEV-Paste strategy (Zhang et al. 2023a) is also used to alleviate overfitting during the long training process. For the ablation study, we use ResNet50 (He et al. 2016) as the image backbone and the image size is downsampled to 256$\times$704. The models are trained for 24 epochs without the CBGS strategy.
4.3 Main Results
We compare GeoBEV with previous state-of-the-art multi-view 3D detectors on the nuScenes val and test set. The experiment results in Tab. 1 show that GeoBEV achieves the best detection accuracy on nuScenes val set. When detecting from images in 256$\times$704 and using ResNet50 as the backbone, GeoBEV outperforms SA-BEV (Zhang et al. 2023a), the previous state-of-the-art, by a significant margin of 2.8% mAP and 2.3% NDS. When increasing image resolution to 512$\times$1408 and using ResNet101 as the backbone, GeoBEV stays ahead of the curve and outperforms SA-BEV by 3.8% mAP and 3.3% NDS.
The results of experiments made on the nuScenes test set are shown in Tab. 2. It can be found that GeoBEV also gets the best performance and outperforms SA-BEV by 1.0% mAP and 1.1% NDS. Those persuasive experiment results highlight the effectiveness of GeoBEV.
We also conduct experiments to figure out whether the BEV representations with high-quality geometric information can combine with the long-term temporal fusion strategy and show the results in Table 3. When using 7 past frames with an interval of 0.5 second to supplement the current scenes, GeoBEV exceeds state-of-the-art long-term temporal methods and outperforms StreamPETR (Wang et al. 2023) by 2.9% mAP and 2.5% NDS.
4.4 Ablation Study
Component Analysis
We evaluate the contributions of our proposed components and show the results in Tab. 4. It can be found that both RC-Sampling and In-Box Label effectively increase the detection accuracy. RC-Sampling increases 2.6% mAP and 3.3% NDS compared to the baseline. Applying In-Box Label and CAI Loss also boosts the performance by 2.2% mAP and 2.2% NDS. After combining the two components, the performance is increased by 4.4% mAP and 4.4% NDS in total.
Baseline RC-Sampling In-Box mAP NDS
BEVDepth 0.337 0.456
0.363 0.489
0.359 0.478
0.381 0.500
Table 4: Ablation study of proposed components. “RC-Sampling” denotes Radial-Cartesian BEV Sampling and “In-Box” denotes the combination of In-Box Label and Centroid-Aware Inner Loss.
Method BEV Size DS mAP NDS FPS
BEVPoolv2 128$\times$128 16 0.337 0.456 22.7
256$\times$256 16 0.344 0.474 16.6
DFA3D 128$\times$128 16 0.335 0.455 20.2
256$\times$256 16 0.344 0.469 11.7
RC-Sampling 128$\times$128 16 0.344 0.465 24.8
256$\times$256 16 0.358 0.482 17.4
256$\times$256 8 0.363 0.489 17.0
Table 5: Ablation study of Radial-Cartesian BEV Sampling. “DS” denotes the downsample factor from the images to the depth scores. “FPS” is the FPS of the whole detector.
Radial-Cartesian BEV Sampling
To evaluate the capacity of RC-Sampling, we compare it with the most efficient LSS-based and Transformer-based feature transformation methods. BEVPoolv2 (Huang and Huang 2022b) and DFA3D (Li et al. 2023a) are chosen as the representatives respectively. From the experiment results in Tab. 5, it can be found the detection accuracy of RC-Sampling outperforms both BEVPoolv2 and DFA3D in generating BEV representations with different resolutions. Besides, RC-Sampling exhibits better real-time performance and achieves the best FPS under the same configuration as other methods. The resolution of depth scores also influences the geometric quality of BEV representations, so we upsample the size of depth scores by utilizing several convolutions to provide fine-grained geometry information to RC-Sampling, which further increases the performance by 0.5% mAP and 0.7% NDS without significantly affects its efficiency.
Label Sigmoid Focal CAI mAP NDS
LiDAR 0.337 0.456
Vanilla In-Box 0.345 0.464
0.347 0.466
0.351 0.470
In-Box 0.356 0.474
0.359 0.478
Table 6: Ablation study of In-Box Label. “Sigmoid” denotes using sigmoid as the activation function. “Focal” denotes using the focal loss while “CAI” denotes using the Centroid-Aware Inner Loss.
In-Box Label
We conduct experiments to evaluate different configurations when applying the In-Box Label as in Tab. 6. When simply replacing the LiDAR label with the vanilla In-Box Label, the performance is increased by 0.8% mAP and 0.8% NDS. It is further improved by 0.2% mAP / 0.2% NDS and 0.4% mAP / 0.4% NDS after using sigmoid as the activation function and letting the depth scores supervised by focal loss. We also compare the performance between vanilla In-Box Label and complete In-Box Label, the results show that the In-Box Label is more in line with the real world and has an advantage of 0.5% mAP and 0.4% NDS. When replacing Focal Loss with Centroid-Aware Inner Loss, there is another 0.3% mAP and 0.4% NDS improvement, which illustrates that inner geometric structure is helpful for the detection.
5 Conclusion
In this paper, we propose a novel multi-view 3D object detector, namely GeoBEV, which generates BEV representation that restores authentic geometric information of the scene. The Radial-Cartesian BEV Sampling simply does high-dimensional matrix multiplication between transposed image features and depth scores to obtain Radial BEV features, which are then transformed into Cartesian BEV features by bilinear sampling. This approach can rapidly generate high-resolution BEV representations while effectively avoiding the presence of vacant feature values. Based on the physics of the real world, In-Box Label can reflect the actual geometric structure of objects, effectively improving the accuracy of the information carried by BEV representation. Centroid-Aware Inner Loss cooperates with In-Box Label to make full of its advantage and also encourages the network to learn the inner geometry of objects.
We conduct extensive experiments on nuScenes dataset and GeoBEV reaches a new state-of-the-art, highlighting the effectiveness of Radial-Cartesian BEV Sampling and In-Box Label. Theoretically, these components can be easily integrated into many existing BEV-based detectors, improving the geometry quality of BEV representation and increasing the detection performance.